Constants Part A A certain simple pendulum has a period on the earth of 1.40 s . What is its period on the surface of Mars, where g-3.71 m/s Request Answer Provide Feedback
Answer #1

Solution,

Time, T = 2 pi sqrt(L/g)

on Earth, L = T^2 g/4 pi^2 = (1.4^2 x 9.8)/(4 pi^2) = 0.4865 m

on Mars, T = 2 pi sqrt(L/g) = 2 pi sqrt(0.4865/3.71) = 2.28 sec

Comment in case any doubt please rate my answer....